OpenStudy (kawaiiiimya):

Can someone help me and explain how you sloved 1. Given e x =6.308 ex=6.308 Round to the nearest thousandths 2. Given e x+1 =4.569 ex+1=4.569 Round to the nearest thousandths 3. Given e 2x−1 =6.897 e2x−1=6.897 Round to the nearest thousandths

OpenStudy (mathmate):

The question posted appears to have a problem of formatting. I assume that `Given e x =6.308 ` is meant to be Given \(e^x =6.308 \) and that you would like to solve for x. Here's an example: Given \(e^x =5.472 \) take \(\log_e\) on both sides, \(\log_e e^x = \log_e 5.472~\Rightarrow ~ x \log_e e =1.69964\Rightarrow x=1.69964\) since \(\log_e e=1\) by definition of logarithm.